Please help me solve this logarithm. The 9 is over the x and their is parenthesis around them.
log3(9)
(x)



Answer by stanbon(75887)

The 9 is over the x and their is parenthesis around them.
y = log3(9/x)
--------------------
y = log3(9)-log3(x)
------
y = 2-log3(x)
=================
Cheers,
Stan H.